| Problem Reports |
|---|
| PR Number | Times | Time Lost | Description | Response |
|---|
| 3968 | 05:00-05:00 | 00:00 | HEFI light source CC error could not be reset, possibly in need of a hardware fix. Unable to measure stack. | |
| 3969 | 05:00-05:00 | 00:00 | Because Jim doesn't hear the dome as often, and because it sounded strange to him, I wanted to report it in case he's picking up on sounds that I've become attuned to because I hear the dome more often. | |
| 3970 | 05:00-05:00 | 00:00 | When Jim was watching the structure air pressure during moves, he noted that it was lower than it has been in the past during moves. Additionally the air bearings sounded fairly squeaky during moves. Wondering if the two are related.
Update: on final move of night, noticed that the structure cut off mid-lift briefly, due to an air pressure fault, before the structure air pressure could catch up. That has not happened in a long time. | |